OpenStudy (anonymous):

Milan will rent a car for the weekend. He can choose one of two plans. The first plan has an initial fee of $59.98 and costs an additional $0.08 per mile driven. The second plan has an initial fee of $49.98 and costs an additional $0.12 per mile driven. How many miles would Milan need to drive for the two plans to cost the same?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@Maliboo

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@kensclark17

OpenStudy (anonymous):

@HeyWassup345

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

Can you make an equation out of this: The first plan has an initial fee of $59.98 and costs an additional $0.08 per mile driven. And can you make another equation out of this? The second plan has an initial fee of $49.98 and costs an additional $0.12 per mile driven.

OpenStudy (anonymous):

Thats the thing I don't know how to do that, Even if you just explained how you know to put what where and why I would have a better understanding

TheSmartOne (thesmartone):

"The first plan has an initial fee of $59.98 and costs an additional $0.08 per mile driven. " We are trying to make an equation and it will be in the form of \(\tt y=mx+b\) So what number is the constant in here? And what is the dependent value that will change based on what x is?

OpenStudy (anonymous):

what does constant mean?

TheSmartOne (thesmartone):

Constant means it is not dependent on any other value. For example, 64 mph is dependent on how many hours you are driving for.

OpenStudy (anonymous):

so 59.98 is the constant

TheSmartOne (thesmartone):

correct

TheSmartOne (thesmartone):

So, now we have: \(\sf y=mx+59.98\) What is the value of m?

OpenStudy (anonymous):

0.80

TheSmartOne (thesmartone):

We can say that x will be how many ever miles are driven. :)

TheSmartOne (thesmartone):

Nope, that is not correct. Look at your numbers carefully again.

OpenStudy (anonymous):

so 0.80=mx+59.98?

TheSmartOne (thesmartone):

.8 is not the correct number. That is not the number they gave you. :P

OpenStudy (anonymous):

I'm not understanding you

TheSmartOne (thesmartone):

"The first plan has an initial fee of $59.98 and costs an additional $\(\bf 0.08\) per mile driven."

OpenStudy (anonymous):

opps i mixed the 8 & 0 sorry haha

OpenStudy (anonymous):

so 0.08=mx+59.98 ?

TheSmartOne (thesmartone):

No.... \(\sf y=.08x+59.98\)

TheSmartOne (thesmartone):

Can you make the other equation based on what you were given now?

OpenStudy (anonymous):

I asked that before and you said no lol

OpenStudy (anonymous):

yes so the second equation should be y=0.12x+49.98

TheSmartOne (thesmartone):

I said no because you had .8 not .08

OpenStudy (anonymous):

Sorry :p

TheSmartOne (thesmartone):

And your second equation is correct :)

OpenStudy (anonymous):

ok so now how do i make them equal or the same or whatever its asking?

TheSmartOne (thesmartone):

Do you know how to solve 2 equations?

OpenStudy (anonymous):

yes, but im solving for y right?

TheSmartOne (thesmartone):

"How many \(\bf miles\) would Milan need to drive for the two plans to cost the same?" We are solving for miles, which is x. y is the overall price at the end of the ride. :)

OpenStudy (anonymous):

ok so the 1st one should be x=-749.75+12.5y & x=-416.5+8.3y?

TheSmartOne (thesmartone):

umm no.

OpenStudy (anonymous):

:(

TheSmartOne (thesmartone):

\(\sf y=.08x+59.98\) \(\sf y=0.12x+49.98\) Do you know how to solve a system of equations by subsituition, elimination, or graphing?

OpenStudy (anonymous):

not really no, since i started home schooling, not having a teacher physically teaching me has caused me to completely lose all math skills i once had. :(

TheSmartOne (thesmartone):

Graphing is the easiest method. Would you want to go with that method to get the answer faster? =P

OpenStudy (anonymous):

yes please !

TheSmartOne (thesmartone):

Here is the graph: https://www.desmos.com/calculator/vndao76k1l

TheSmartOne (thesmartone):

At what point do the two lines intersect at?

TheSmartOne (thesmartone):

The point should be \(\sf (x,y)\) And your answer will be \(\sf x\)

OpenStudy (anonymous):

I have no idea. :( I'm sorry